# Enhancing LLM Language Adaption through Cross-lingual In-Context Pre-training

### v1

@@ Abstract
Linjuan Wu1111Work done during internship at Tongyi Lab., Haoran Wei2222Contributed equally., Huan Lin2, Tianhao Li2, Baosong Yang2, Weiming Lu1, [^3] Corresponding authors. 1Zhejiang University 2Tongyi Lab, Alibaba Group 1{wulinjuan525,luwm}@zju.edu.cn 2{funan.whr, lilai.lh, chongsheng.lth, yangbaosong.ybs}@alibaba-inc.com
@@ 2 Related Work
Figure 2: The implementation process of our method, CrossIC-PT, which constructs cross-lingual in-contexts based on Wikipedia data and performs continued pre-training (CPT) on existing multilingual models. Here, $N$ represents the input window length of the model. The $T$ indicates the title of the articles, and $L$ indicates the target language.
Cross-lingual supervision signals have been proven effective in enhancing LLMs’ cross-lingual transfer abilities Singh et al. (2024); Yamada and Ri (2024). Most methods rely on bilingual corpora as explicit supervision signals Zhang et al. (2024b); Ming et al. (2024); Ji et al. (2024); Gosal et al. (2024); Gilabert et al. (2024). Some works, like Zhang et al. (2024b) distills translation pairs from LLMs through back-translation to create supervision signals. Others, such as Singh et al. (2024); Yamada and Ri (2024), apply code-switching techniques to replace or augment words with English translations.  Yoo et al. (2024) also explores code-switching at various levels using curriculum learning. However, parallel corpora have restricted types, domains (most bilingual corpora are short sentence level bitexts, and usually extracted from news websites), and quantity. Synthetic parallel documents built by back-translation, however, are limited in text Quality. In contrast, our method constructs semantically related document pairs from the authentic data on the Internet, which is more scalable and less problematic.
@@ 3 Method
Multilingual LLM pre-training typically randomly packs documents from different languages into the fixed-size context window. We hypothesize that concatenating semantically related documents from English and a target language enhances cross-lingual transfer. This approach allows the model to predict the next token using both monolingual and cross-lingual context within the concatenated sequence. We call this concatenated sample Cross-lingual In-context data, where English serves as the guiding context for learning the target language. Based on this, we propose CrossIC-PT, a pre-training method leveraging cross-lingual in-context data.
@@ 3.4 Data Augmentation via Retrieval
To validate our approach, we use the Wikipedia corpus, which includes data in nearly 200 languages linked by matched titles. While the content across languages is not strictly parallel, it covers the same topics, making it suitable for our needs. To enhance the generalization of our method, we introduce a cross-lingual semantic retrieval framework based on the FAISS similarity search tool Johnson et al. (2019), as shown in Fig.3. This framework augments the training data by incorporating relevant English articles from the Fineweb_edu Lozhkov et al. (2024) dataset, retrieved using title and content keywords (up to 10 per article) extracted from the Wikipedia data.
@@ 4.1 Training Data
Our training data is primarily sourced from Wikipedia (denoted as W), with token counts for English and each target language listed in Table 1. We selected six target languages $L$: Arabic (ar), Spanish (es), Japanese (ja), Korean (ko), Portuguese (pt), and Thai (th). To further expand the dataset, we retrieved relevant English data from a subset of Fineweb_edu (denoted as F), which has a file size of 17.44GB. The token counts for the augmented data are also provided in Table 1.
@@
Llama-2-7B base 24.77 37.10 37.76 35.05 40.90 23.27
EMMA-500 30.14 31.18 32.77 32.49 28.06 33.31
Llama-3.1-8B base 37.96 42.11 43.02 43.82 44.36 38.79
LEIA 37.04±0.49 44.03±0.24 44.86±0.89 44.11±0.48 44.48±0.58 42.90±0.82
Mix-PT 38.09 43.46 44.81 44.75 46.45 42.38
CrossIC-PT 40.57 45.49 47.27 46.87 49.09 43.51
Qwen2.5-7B base 50.91 54.71 56.95 55.52 56.49 53.81
Mix-PT 54.48 58.71 57.69 57.39 60.30 56.19
CrossIC-PT 55.97 59.44 59.00 59.03 61.59 57.33
Qwen2.5-1.5B base 37.83 43.90 42.26 39.75 44.35 41.40
Mix-PT 38.14 44.37 41.85 39.48 45.63 40.92
CrossIC-PT 40.21 45.09 43.96 41.47 48.25 42.23
Table 2: The average results of our CrossIC-PT model, based on three base LLMs (Llama-3.1-8B, Qwen2.5-7B, and Qwen2.5-1.5B), are compared with corresponding baselines across six target languages. The cross-lingual in-context datasets used in CrossIC-PT are sourced from Wikipedia.
@@ 4.3 Benchmark
We evaluated our models on several tasks from the latest multilingual and multitask benchmark, P-MMEVAL Zhang et al. (2024a), which includes: generation (FLORES-200 Costa-jussà et al. (2022)), understanding (XNLI Conneau et al. (2018), MHELLASWAG ***https://huggingface.co/datasets/alexandrainst/m_hellaswag), knowledge (MMMLUhttps://huggingface.co/datasets/openai/MMMLU), logical reasoning (MLOGIQA), and mathematical reasoning (MGSM Shi et al. (2023)). To further assess the models’ paragraph comprehension abilities, we incorporated a reading comprehension task (MRC). The MRC test data includes TydiQA-GoldP Clark et al. (2020) for Arabic (ar) and Korean (ko), XQuAD Artetxe et al. (2020) for Spanish (es), Portuguese (pt), and Thai (th), and 1,200 samples from JaQuAD So et al. (2022) for Japanese (ja). Details of the evaluation setting can be found in Appendix A.
@@ 4.4 Baselines
In addition to the base models (Llama-3.1-8B, Qwen2.5-7B, and Qwen2.5-1.5B), we included the following baselines:
  • EMMA-500 Ji et al. (2024): A model CPT on Llama-2-7B Touvron et al. (2023) with 136B tokens covering over 500 languages.
  • LEIA Yamada and Ri (2024): A method that randomly adds English translations of entities to target-language Wikipedia data for pre-training, leveraging cross-lingual entity supervision. We reproduced this method using the provided code to construct the data and perform CPT on Llama-3.1-8B, ensuring the target-language token count matched ours. We conducted experiments with three random seeds (32, 111, 222) and reported the mean and variance of the results.
  • Mix-PT: A method that uses our title-matched article pairs from Fig.2(a) for pre-training.
@@ 4.5.1 Base Results
The average results of the baselines and our method, based on data from Wikipedia in six languages, are shown in Table 2. Detailed results for each task can be found in Appendix B. CrossIC-PT consistently improves the performance of the base LLMs and outperforms other baselines, demonstrating the effectiveness of using semantically related cross-lingual in-context corpora for pre-training.
@@
Compared to the base LLMs, our CrossIC-PT method improves performance by 3.79%, 3.99%, and 1.95% on Llama-3.1-8B, Qwen2.5-7B, and Qwen2.5-1.5B, respectively, across six languages. Notably, in Portuguese (pt), CrossIC-PT improves performance by 4.73% on Llama-3.1-8B, surpassing the strongest baseline by 2.64%. The performance gains for Qwen2.5 models are more pronounced as model size increases, which may be attributed to the fact that CPT performance is influenced by the initial capabilities of the model.
@@
Data Model Languages
ar es ja ko pt th
Llama-3.1-8B 37.96 42.11 43.02 44.14 44.36 38.79
W Mix-PT 38.09 43.46 44.81 44.75 46.45 42.38
CrossIC-PT 40.57 45.49 47.27 46.87 49.09 43.51
W+F Mix-PT 40.19 44.58 44.75 44.48 46.62 42.05
CrossIC-PT 41.18 46.93 48.10 47.32 49.97 43.72
Table 3: The results of our CrossIC-PT model and Mix-PT baseline with Wikipedia-based data and augmented data constructed on Wikipedia and Fineweb_edu data.
@@ 4.5.2 Results of Data Augmentation
To explore the generalization of our method, we propose a cross-lingual semantic retrieval framework (shown in Fig. 3) to augment the training data, with results reported in Table 3. After retrieval, the data volume increased by 0.06B–0.23B. Although this is a relatively small increase, it improved the average performance of our method by 0.73%. This demonstrates that even when English data is not perfectly aligned with target languages, semantically related still aids cross-lingual transfer. The simplicity of the semantic similarity retrieval process allows easy extension to various data sources.
@@
Figure 5: Ablation results of CrossIC-PT without optimized sliding window mechanism (Opt-SWM).
@@
4.6 Ablation Study on Sliding Window Mechanism
We conduct an ablation study to assess the impact of our optimized sliding window mechanism (Opt-SWM), which introduces the [SPLIT] token and ensures each window starts after the last [SPLIT] token. Specifically, we compare the performance of CrossIC-PT with and without Opt-SWM, denoted as CrossIC-PT w/o Opt-SWM. The results shown in Fig.5 reveal that even without Opt-SWM, using only window-split cross-lingual in-context data, CrossIC-PT consistently improves performance across all languages. The addition of the optimized sliding window mechanism further enhances performance, highlighting its role in maintaining cross-lingual in-context coherence and improving language transfer. This demonstrates the effectiveness of all the steps in our design.
@@ 5 Analysis
We believe concatenating semantically related English and target language text in cross-lingual in-context data helps the model better understand the target language guided by the English context. Thus, we set the order as English first, followed by the target language. To verify if this direction is more beneficial, we analyze the concatenation order and test the model’s performance in English to ensure there is no catastrophic forgetting.
5.1 Analysis of Concatenation Direction
@@ 5.2 Performance on English Tasks
To prevent catastrophic forgetting, it’s important to ensure English performance is maintained. To verify this, we tested the performance of six target language models on English tasks, using the same tasks as before. The results are shown in Fig.6.
The upper part of Fig.6 shows the average performance of each target language model on English tasks, with the x-axis ordered by the performance gap between Llama-3.1-8B’s performance on the target language and English. The trend suggests that a larger performance gap corresponds to a greater impact on English performance after training. For example, the English performance of Thai (th) and Arabic (ar) is lower. However, it is primarily due to a significant drop in one task. To further investigate, the lower part of Fig.6 presents the statistical significance ("p") of the performance differences between target language models and the base model, Llama-3.1-8B, across seven tasks. The results show that, except for the Thai model trained with data augmentation (which exhibits a significant drop in English performance), there are no significant differences for other target language models. This suggests that CrossIC-PT improves performance in target languages while effectively preserving English capabilities. We believe this is likely due to the inclusion of at least 50% of English tokens in the cross-lingual in-context corpus, which helps mitigate severe forgetting. This result further validates the robustness and practicality of CrossIC-PT for cross-lingual transfer.
@@ Acknowledgements
This work was supported by the Alibaba Research Intern Program.
@@ Appendix A The Setting for Evaluation
Table 5: Task Prompts."[]" represents optional content. For the FLORESE task, the "[source]" indicates the source language, and "[target]" indicates the target language of translation. For MGSM, "[The answer is ]" is the translation of "The answer is " according to the test language.
Appendix B Results of Tasks

### v2

@@ Abstract
Linjuan Wu1111Work done during internship at Tongyi Lab., Haoran Wei2222Contributed equally., Huan Lin2, Tianhao Li2, Baosong Yang2, Fei Huang2444Google Scholar ID is 9r98PpoAAAAJ, Weiming Lu1, [^3] Corresponding authors. 1Zhejiang University 2Tongyi Lab, Alibaba Group 1{wulinjuan525,luwm}@zju.edu.cn 2{funan.whr, lilai.lh, chongsheng.lth, yangbaosong.ybs, f.huang}@alibaba-inc.com
@@ 2 Related Work
Figure 2: The implementation process of our method, CrossIC-PT, which constructs cross-lingual in-contexts based on Wikipedia data and performs continued pre-training (CPT) on existing multilingual models. Here, NN represents the input window length of the model. The $T$ indicates the title of the articles, and $L$ indicates the target language.
Cross-lingual supervision signals have been proven effective in enhancing LLMs’ cross-lingual transfer abilities Singh et al. (2024); Yamada and Ri (2024); PLUG. Most methods rely on bilingual corpora as explicit supervision signals Zhang et al. (2024b); Ming et al. (2024); Ji et al. (2024); Gosal et al. (2024); Gilabert et al. (2024). Some works, like Zhang et al. (2024b) distills translation pairs from LLMs through back-translation to create supervision signals. Others, such as Singh et al. (2024); Yamada and Ri (2024), apply code-switching techniques to replace or augment words with English translations.  Yoo et al. (2024) also explores code-switching at various levels using curriculum learning. However, parallel corpora have restricted types, domains (most bilingual corpora are short sentence level bitexts, and usually extracted from news websites), and quantity. Synthetic parallel documents built by back-translation, however, are limited in text Quality. In contrast, our method constructs semantically related document pairs from the authentic data on the Internet, which is more scalable and less problematic.
@@ 3 Method
Multilingual LLM pre-training typically packs documents from different languages randomly into the fixed-size context window. We hypothesize that concatenating semantically related English and target language corpora, predicting the next words based upon not only monolingual and cross-lingual context could enhance cross-lingual transfer ability. We call this concatenated sample Cross-lingual In-context data, where English serves as the guiding context for learning the target language. Based on this, we propose CrossIC-PT, a pre-training method leveraging cross-lingual in-context data.
@@ 3.4 Data Augmentation via Retrieval
To validate our approach, we use the Wikipedia corpus, which includes data in nearly 200 languages linked by matched titles. While the content across languages is not strictly parallel, it covers the same topics, making it suitable for our needs. To enhance the generalization of our method, we introduce a cross-lingual semantic retrieval framework based on the FAISS similarity search tool Johnson et al. (2019)***We used the embedding from https://huggingface.co/Alibaba-NLP/gte-multilingual-base., as shown in Fig.3. This framework augments the training data by incorporating relevant English articles from the Fineweb_edu Lozhkov et al. (2024) dataset, retrieved using title and content keywords (up to 10 per article) extracted from the Wikipedia data.
@@ 4.1 Training Data
Our training data is primarily sourced from WikipediaWe used the 20240720 wiki-dumps and processed them with wikiextractor (https://github.com/attardi/wikiextractor) to remove boilerplate text and extract article content. (denoted as W), with token counts for English and each target language listed in Table 1. We selected six target languages $L$: Arabic (ar), Spanish (es), Japanese (ja), Korean (ko), Portuguese (pt), and Thai (th). To further expand the dataset, we retrieved relevant English data from a subset of Fineweb_edu (denoted as F), which has a file size of 17.44GB. The token counts for the augmented data are also provided in Table 1.
@@
Llama-3.1-8B base 37.96 42.11 43.02 43.82 44.36 38.79 41.68
LEIA 37.04±0.49 44.03±0.24 44.86±0.89 44.11±0.48 44.48±0.58 42.90±0.82 42.98±0.25
X-MONO-PT 39.11 44.35 43.57 44.24 44.04 42.09 42.90
F Mix-PT 38.24 44.09 44.09 44.12 45.65 41.44 42.94
CrossIC-PT 39.66 44.57 45.61 46.02 47.32 42.30 44.24
W Mix-PT 38.09 43.46 44.81 44.75 46.45 42.38 43.32
CrossIC-PT 40.57 45.49 47.27 46.87 49.09 43.51 45.47
W+F Mix-PT 40.19 44.58 44.75 44.48 46.62 42.05 43.78
CrossIC-PT 41.18 46.93 48.10 47.32 49.97 43.72 46.20
Qwen-2.5-7B base 50.91 54.71 56.95 55.52 56.49 53.81 54.73
Mix-PT 54.48 58.71 57.69 57.39 60.30 56.19 57.46
CrossIC-PT 55.97 59.44 59.00 59.03 61.59 57.33 58.73
Qwen-2.5-1.5B base 37.83 43.90 42.26 39.75 44.35 41.40 41.58
Mix-PT 38.14 44.37 41.85 39.48 45.63 40.92 41.73
CrossIC-PT 40.21 45.09 43.96 41.47 48.25 42.23 43.54
Table 2: The average results of our CrossIC-PT model, based on three base LLMs (Llama-3.1-8B, Qwen2.5-7B, and Qwen2.5-1.5B), are compared with corresponding baselines across six target languages. The cross-lingual in-context datasets used in methods based on Qwen2.5-7B and Qwen2.5-1.5B are sourced from Wikipedia(W).
@@ 4.3 Benchmark
We evaluated our models on several tasks from the latest multilingual and multitask benchmark, P-MMEVAL Zhang et al. (2024a), which includes: generation (FLORES-200 Costa-jussà et al. (2022)), understanding (XNLI Conneau et al. (2018), MHELLASWAG https://huggingface.co/datasets/alexandrainst/m_hellaswag), knowledge (MMMLU§§§https://huggingface.co/datasets/openai/MMMLU), logical reasoning (MLOGIQA), and mathematical reasoning (MGSM Shi et al. (2023)). To further assess the models’ paragraph comprehension abilities, we incorporated a reading comprehension task (MRC). The MRC test data includes TydiQA-GoldP Clark et al. (2020) for Arabic (ar) and Korean (ko), XQuAD Artetxe et al. (2020) for Spanish (es), Portuguese (pt), and Thai (th), and 1,200 samples from JaQuAD So et al. (2022) for Japanese (ja). Details of the evaluation setting can be found in Appendix A.
@@ 4.4 Baselines
In addition to the base models (Llama-3.1-8B, Qwen2.5-7B, and Qwen2.5-1.5B), we included the most relevant baseline Mix-PT, which uses title-matched article pairs (Fig. 2(a)) for pre-training but without cross-lingual text concatenation.
Based on Llama-3.1-8B we also included two other baselines:(1)LEIA Yamada and Ri (2024): A method that randomly adds English translations of entities to target-language Wikipedia data for pre-training, leveraging cross-lingual entity supervision. We reproduced this method using the provided code to construct the data and perform CPT on Llama-3.1-8B, ensuring the target-language token count matched ours. We conducted experiments with three random seeds (32, 111, 222) and reported the mean and variance of the results. (2)X-MONO-PT: A method that uses the target language data from our title-matched article pairs in Fig.2(a) for pre-training.
@@ 4.5.1 Base Results
ALl average six languages results of the baselines and our method, based on different training data volume from Wikipedia (W) and Fineweb_edu (F) are shown in Table 2. Detailed results for each languages can be found in Appendix B. CrossIC-PT consistently improves the performance of the base LLMs and outperforms other baselines, demonstrating the effectiveness of using semantically related cross-lingual in-context corpora for pre-training.
@@
Compared to the base LLMs, our CrossIC-PT method trained with only Wikipedia(W) data improves performance by 3.79%, 3.99%, and 1.95% on Llama-3.1-8B, Qwen2.5-7B, and Qwen2.5-1.5B, respectively, across six languages. Notably, in Portuguese (pt), CrossIC-PT improves performance by 4.73% on Llama-3.1-8B, surpassing the strongest baseline by 2.64%. The performance gains for Qwen2.5 models are more pronounced as model size increases, which may be attributed to the fact that CPT performance is influenced by the initial capabilities of the model.
@@ 4.5.2 Results of Data Augmentation
To explore the generalization of our method, we propose a cross-lingual semantic retrieval framework (Fig. 3) to augment the training data by expanding Wikipedia (W) with FineWeb-edu (F). After retrieval, the data volume increased by 0.06B–0.23B tokens. Although this is a relatively small increase, it improved the average performance of our method by 0.73%. Even when using only the augmented data (F), our method still achieved a 1.26% improvement over the strong baseline LEIA. This demonstrates that even if English data is not perfectly aligned with target languages, semantic similarity still facilitates cross-lingual transfer. Moreover, the simplicity of our retrieval process allows easy extension to various data sources.
@@
Figure 5: Ablation results of CrossIC-PT.
@@
4.6 Ablation Study
We conduct ablation studies to evaluate two key components of our approach: (1) the optimized sliding window mechanism (Opt-SWM) and (2) the semantic related of cross-lingual contexts. First, comparing CrossIC-PT with and without Opt-SWM (denoted as CrossIC-PT w/o Opt-SWM), Fig.5 shows that while window-split cross-lingual contexts alone improve performance across all languages, adding Opt-SWM further enhances results by maintaining context coherence. Second, when replacing semantically similar pairs with randomly paired bilingual documents (CrossIC-PT_random), performance degrades significantly - though still surpassing Mix-PT baselines - confirming the importance of our semantic similarity strategy. These results collectively validate the effectiveness of each design component in CrossIC-PT.
@@ 5 Analysis
5.1 Concatenation Direction
We believe that placing semantically related English text before target-language content helps models better learn from cross-lingual contexts. Thus, we set the order as English first, followed by the target language. To verify if this direction is more beneficial, we analyze the concatenation order.
@@ 5.2 Performance on English Tasks
To prevent catastrophic forgetting during continual pre-training, it’s important to ensure English performance is maintained. To verify this, we tested the performance of six target language models on English tasks, using the same tasks as before. The results are shown in Fig.6.
The upper part of Fig.6 shows the average performance of each target language model on English tasks, with the x-axis ordered by the performance gap between Llama-3.1-8B’s performance on the target language and English. The trend suggests that a larger performance gap corresponds to a greater impact on English performance after training. For example, the English performance of Thai (th) and Arabic (ar) is lower. However, it is primarily due to a significant drop in one task. To further investigate, the lower part of Fig.6 presents the statistical significance (”p”) of the performance differences between target language models and the base model, Llama-3.1-8B, across seven tasks. The results show that, except for the Thai model trained with data augmentation (which exhibits a significant drop in English performance), there are no significant differences for other target language models. This suggests that CrossIC-PT improves performance in target languages while effectively preserving English capabilities. We believe this is likely due to the inclusion of at least 50% of English tokens in the cross-lingual in-context corpus, which helps mitigate severe forgetting. This result further validates the robustness and practicality of CrossIC-PT for cross-lingual transfer.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (No. 62376245), the Key Research and Development Program of Zhejiang Province, China (No. 2024C01034), the Fundamental Research Funds for the Central Universities (226-2024-00170), National Key Research and Development Project of China (No. 2018AAA0101900), and MOE Engineering Research Center of Digital Library, and the Alibaba Research Intern Program.
@@ Appendix A The Setting for Evaluation
Table 4: Task Prompts.”[]” represents optional content. For the FLORESE task, the ”[source]” indicates the source language, and ”[target]” indicates the target language of translation. For MGSM, ”[The answer is ]” is the translation of ”The answer is ” according to the test language.
For the FLORES-200 generation tasks, we employed SacreBLEUhttps://github.com/mjpost/sacreBLEU with the default configuration: nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.0.0. For languages lacking whitespace-based word boundaries (Japanese [ja], Korean [ko], and Thai [th]), we introduced a pre-tokenization step prior to BLEU computation, implemented as follows:
1class NonASCIITokenizer(object):
2 def __init__(self):
3 self.is_cjk = re.compile(
[…]
8 ")"
9 )
10
11 def __call__(self, sent):
12 sent = sent.strip()
13 chs = list(sent)
14 line_chtok = []
15 for ch in chs:
16 if self.is_cjk.match(ch):
17 line_chtok.append()
18 line_chtok.append(ch)
19 line_chtok.append()
20 else:
21 line_chtok.append(ch)
22 line_chtok = trim_multiple_space(line_chtok)
23 return .join(line_chtok)
Appendix B Results of Per-Languages